OpenStudy (anonymous):

What is the missing exponent? (106) ?= 10 -18

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

please use `^` for exponents like 2 to the 3rd power = 2^3 \[\huge\rm (10^6)^{??}=10^{-18}\]

Nnesha (nnesha):

sorry last one was wrong

Nnesha (nnesha):

now apply the exponent rule \[\huge\rm (x^m )^n =x^{m \times n}\]